OpenStudy (anonymous):

I'm doing a review sheet right now and I really need the help please. The first term of an arithmetic sequence is 2, the fourth term is 11. Find the sum of the first 50 terms.

OpenStudy (anonymous):

Listing down what we have: \[a =2 \] \[a_{4}= a + (n-1)d =11\] Solving for d: \[(2)+3d=11\] \[3d=9\] \[d=3\] Sum of the first 50 terms: \[\frac{ 50}{ 2 }(2a + 49d)\] Since you are given "a" and have found "d", sub them in to solve the problem.